# FedP2EFT: Federated Learning to Personalize Parameter Efficient Fine-Tuning for Multilingual LLMs

### v1

@@ 4.1.3 Baselines
For each baseline, we either follow best practices recommended by the corresponding authors or employ a simple grid search and pick the best performing hyperparameters. Full details in Appendix. A.
@@ 4.2 Results on Text Classification
We evaluate our approach in a typical FL setup, where the pretrained model is first trained using Standard FL with full fine-tuning and the resulting base model is then personalized to each client. Tables 1 & 2 show the mean and standard deviation (SD) of the accuracy for each language in our MasakhaNEWS setup for seen and unseen pool respectively (similarly for XNLI in Appendix Tables 8 & 9). In addition, we also show the number of languages, labelled “Wins”, an approach is best performing for each budget $r$.
@@
The results in all four tables show that federated learning to personalize methods (FedL2P and FedP2EFT) outperform the other baselines in most cases. Non-FL rank selection approaches (AdaLoRA and BayesTune-LoRA), on the other hand, tend to overfit and/or struggle to learn an optimal rank structure given the limited number of samples in each client. Comparing FedL2P and FedP2EFT, FedP2EFT largely surpass FedL2P with a few exceptions, indicating that learning to personalize LoRA rank structure is the better hyperparameter choice than personalizing learning rates; this finding is also aligned with recent LLM-based optimizer findings (Zhao et al., 2025), which shows that Adam’s performance is robust with respect to its learning rate.
4.2.1 FedP2EFT’s Complementability with Personalized FL Works.
$\mathbf{r}$ Approach bg hi es el vi tr de ur en zh th sw ar fr ru Wins
2 LoRA 45.80±0.28 42.80±0.16 48.73±0.25 50.87±0.09 53.00±0.00 48.00±0.33 49.87±0.09 41.93±0.09 46.53±0.34 44.40±0.16 42.53±0.25 51.80±0.16 46.93±0.25 48.07±0.19 50.53±0.25 0
AdaLoRA 44.07±0.09 41.20±0.00 47.47±0.09 50.00±0.00 52.40±0.00 46.53±0.09 48.00±0.00 38.80±0.00 44.67±0.09 42.20±0.00 40.67±0.09 50.40±0.00 45.00±0.00 46.00±0.00 48.80±0.00 0
BayesTune-LoRA 43.80±0.00 40.60±0.00 47.20±0.00 50.00±0.00 52.40±0.00 46.20±0.00 47.40±0.16 38.80±0.00 44.27±0.09 41.60±0.00 40.53±0.09 49.80±0.00 44.73±0.09 46.00±0.00 48.07±0.19 0
FedL2P 47.47±2.78 44.53±3.18 50.27±2.94 51.47±1.39 53.60±0.71 50.07±2.81 50.93±2.62 44.80±4.53 49.07±4.35 46.53±4.01 44.40±3.68 52.27±1.27 48.40±2.97 49.47±2.96 51.60±2.27 0
FedP2EFT 64.40±0.16 57.80±1.34 58.53±1.32 59.73±2.58 60.80±0.71 58.87±2.87 55.00±0.16 63.67±0.19 55.93±0.19 56.27±0.34 58.33±0.25 59.47±0.47 55.20±0.16 57.53±0.19 55.20±0.43 15
4 LoRA 48.73±0.38 46.07±0.66 53.27±0.09 52.53±0.09 53.33±0.09 50.27±0.25 52.80±0.28 46.13±0.25 51.67±0.25 48.53±0.19 45.47±0.19 53.33±0.09 50.47±0.41 50.27±0.09 52.87±0.19 0
AdaLoRA 44.00±0.00 41.13±0.09 47.40±0.00 50.00±0.00 52.40±0.00 46.40±0.00 47.67±0.09 38.80±0.00 44.47±0.09 41.93±0.09 40.60±0.00 50.27±0.09 44.87±0.09 46.00±0.00 48.80±0.00 0
BayesTune-LoRA 44.00±0.00 41.00±0.00 47.20±0.00 50.00±0.00 52.40±0.00 46.33±0.09 47.67±0.09 38.80±0.00 44.47±0.09 41.80±0.00 40.60±0.00 50.00±0.00 44.80±0.00 46.00±0.00 48.67±0.09 0
FedL2P 48.67±2.36 45.20±2.26 52.13±1.20 51.87±0.52 53.87±0.34 50.27±0.84 52.00±0.99 47.00±3.69 50.73±2.59 47.27±1.95 45.07±1.65 52.67±0.52 49.67±1.59 51.20±2.26 52.47±1.37 0
FedP2EFT 64.47±1.18 64.00±0.98 64.40±0.91 63.07±0.41 64.13±0.34 64.87±0.68 63.33±1.67 64.47±0.62 56.60±0.28 63.93±1.16 62.60±1.73 65.00±0.33 64.13±0.81 61.73±1.23 61.80±0.65 15
8 LoRA 55.80±0.16 51.73±0.25 55.73±0.09 55.07±0.25 54.40±0.00 52.80±0.16 54.60±0.00 57.47±0.09 55.53±0.09 54.00±0.33 51.93±0.09 54.13±0.09 53.07±0.09 56.27±0.09 54.60±0.16 0
AdaLoRA 43.80±0.00 40.87±0.09 47.27±0.09 50.00±0.00 52.40±0.00 46.27±0.09 47.73±0.09 38.80±0.00 44.40±0.00 41.87±0.09 40.60±0.00 50.00±0.00 44.80±0.00 46.00±0.00 48.40±0.16 0
BayesTune-LoRA 44.13±0.09 41.20±0.00 47.47±0.09 50.20±0.00 52.40±0.00 46.47±0.09 47.93±0.09 39.13±0.09 44.93±0.09 42.53±0.09 40.87±0.09 50.47±0.09 45.13±0.09 46.13±0.09 48.80±0.00 0
FedL2P 52.20±0.59 48.73±0.25 54.20±0.16 53.13±0.19 54.33±0.09 51.53±0.19 53.60±0.16 51.00±0.86 54.27±0.25 51.73±0.25 47.07±0.52 53.60±0.16 51.87±0.19 52.93±0.34 54.60±0.16 0
FedP2EFT 67.27±0.25 66.60±0.33 68.87±0.41 65.00±0.43 66.73±0.09 67.80±0.28 67.13±0.19 67.60±0.16 64.00±0.99 67.67±0.38 66.53±0.09 67.33±0.25 67.60±0.33 66.53±0.25 69.33±0.19 15
16 LoRA 64.73±0.09 64.27±0.19 65.60±0.28 65.33±0.19 63.80±0.00 64.27±0.09 64.20±0.71 64.80±0.16 57.93±2.46 65.93±0.25 58.33±0.52 65.13±0.19 60.53±0.62 64.53±0.47 61.73±0.74 1
AdaLoRA 43.87±0.09 40.73±0.09 47.20±0.00 50.00±0.00 52.40±0.00 46.33±0.09 47.67±0.09 38.80±0.00 44.27±0.09 41.80±0.00 40.60±0.00 50.00±0.00 44.80±0.00 46.00±0.00 48.27±0.09 0
BayesTune-LoRA 44.60±0.00 41.67±0.19 47.80±0.00 50.20±0.00 52.87±0.09 47.13±0.09 48.33±0.19 40.60±0.00 45.40±0.00 43.60±0.00 41.00±0.00 50.80±0.00 45.33±0.09 46.60±0.00 49.40±0.16 0
FedL2P 54.47±1.86 52.73±0.50 55.80±0.43 54.73±0.47 55.67±1.09 54.80±0.00 54.73±0.25 57.67±2.78 55.67±0.09 54.93±0.41 52.60±2.79 53.93±0.41 53.47±0.25 55.60±2.12 55.20±0.00 0
FedP2EFT 67.87±0.74 67.00±0.43 69.53±0.57 64.93±0.41 67.00±0.33 68.00±0.00 67.20±0.43 68.00±0.59 65.20±1.07 67.60±0.16 66.60±0.57 67.33±0.09 67.80±0.28 66.13±0.41 69.87±0.25 14
$\mathbf{r}$ Approach bg hi es el vi tr de ur en zh th sw ar fr ru Wins
2 LoRA 54.10±0.10 55.80±0.00 57.10±0.10 55.10±0.30 56.30±0.10 54.30±0.10 55.10±0.10 53.60±0.20 61.50±0.10 54.80±0.00 50.70±0.10 52.50±0.10 53.40±0.20 53.10±0.10 55.30±0.10 0
AdaLoRA 53.33±0.09 54.53±0.09 55.33±0.09 52.80±0.28 54.07±0.09 52.87±0.09 53.93±0.09 52.87±0.19 60.40±0.16 53.07±0.09 49.40±0.16 50.67±0.09 51.40±0.16 52.80±0.16 54.20±0.16 0
BayesTune-LoRA 53.40±0.00 54.40±0.00 55.40±0.00 53.10±0.10 54.00±0.20 52.90±0.10 54.10±0.10 53.20±0.00 60.00±0.00 53.20±0.00 49.30±0.10 50.50±0.10 51.50±0.10 52.60±0.20 54.00±0.00 0
FedL2P 64.70±1.10 64.20±2.80 67.90±1.30 68.50±1.10 68.60±0.20 65.20±4.00 67.30±0.70 67.00±0.40 71.50±0.30 64.90±0.50 61.70±1.70 64.60±2.80 65.80±2.00 63.40±0.00 67.40±0.40 0
FedP2EFT 70.67±0.77 70.07±1.04 73.87±1.54 73.47±0.19 73.40±0.59 71.93±2.08 73.27±0.57 71.87±0.62 74.53±0.52 73.80±0.43 65.60±2.97 69.93±1.16 74.40±0.49 71.27±1.52 72.73±1.88 15
4 LoRA 56.67±0.34 59.07±0.25 59.13±0.41 57.33±0.25 59.00±0.28 56.93±0.25 57.40±0.59 55.93±0.09 63.47±0.25 57.27±0.09 52.47±0.34 54.53±0.25 56.40±0.33 55.93±0.09 58.40±0.33 0
AdaLoRA 53.33±0.09 54.40±0.16 55.27±0.09 53.20±0.00 54.13±0.19 52.93±0.09 54.00±0.00 52.87±0.09 60.33±0.25 53.13±0.09 49.20±0.16 50.60±0.16 51.27±0.09 52.60±0.16 54.00±0.00 0
BayesTune-LoRA 53.27±0.09 54.53±0.19 55.40±0.16 53.00±0.16 54.53±0.25 52.87±0.09 54.00±0.00 53.07±0.09 60.60±0.16 53.07±0.09 49.33±0.19 51.07±0.09 51.47±0.09 52.60±0.00 54.53±0.19 0
FedL2P 66.50±1.10 65.40±1.80 69.90±1.90 70.50±0.90 70.10±0.30 66.90±4.10 69.70±0.90 68.20±0.60 72.80±0.20 67.40±1.20 62.60±2.00 65.70±2.90 67.70±1.30 66.50±1.10 69.00±0.40 0
FedP2EFT 71.33±0.34 70.07±1.09 73.27±2.29 73.27±0.68 72.60±0.28 71.87±2.32 74.60±0.16 72.93±0.38 75.00±0.28 74.33±1.16 66.13±3.21 68.13±0.34 75.13±1.65 72.20±0.33 73.47±1.60 15
8 LoRA 60.60±0.28 62.47±0.09 63.73±0.09 62.60±0.33 64.67±0.09 61.93±0.41 62.53±0.09 60.33±0.19 67.33±0.25 61.40±0.16 56.53±0.09 59.27±0.09 61.33±0.25 58.87±0.09 63.13±0.19 0
AdaLoRA 53.33±0.09 54.53±0.19 55.27±0.09 52.80±0.28 54.27±0.25 53.00±0.00 53.93±0.09 52.93±0.09 60.13±0.34 53.00±0.00 49.20±0.16 50.27±0.19 51.20±0.16 52.73±0.38 53.93±0.09 0
BayesTune-LoRA 53.33±0.09 54.40±0.00 55.53±0.19 53.20±0.00 54.67±0.09 53.00±0.00 54.07±0.09 53.13±0.25 60.60±0.00 53.47±0.09 49.47±0.09 51.33±0.34 51.73±0.19 52.87±0.09 54.40±0.16 0
FedL2P 66.40±0.60 65.50±1.10 70.50±3.50 70.10±0.50 70.50±1.70 68.20±2.20 69.60±0.40 67.50±1.30 72.60±0.60 67.10±1.30 61.80±1.00 65.60±1.80 68.20±1.00 67.20±1.40 69.10±1.70 0
FedP2EFT 70.53±0.25 69.27±0.77 73.33±1.79 71.27±1.39 71.33±0.52 70.53±2.10 75.33±0.41 73.07±0.77 74.40±0.75 73.93±1.11 66.73±3.80 68.07±0.94 74.67±1.18 71.87±0.19 72.93±0.96 15
16 LoRA 67.13±0.34 67.80±0.00 69.47±0.09 71.27±0.19 69.20±0.00 68.07±0.38 69.00±0.33 68.73±0.25 71.47±0.25 68.00±0.16 62.80±0.00 67.33±0.25 66.80±0.16 65.07±0.19 67.33±0.09 0
AdaLoRA 53.40±0.00 54.40±0.00 55.40±0.00 52.60±0.16 54.00±0.00 52.93±0.09 53.93±0.09 52.93±0.09 60.27±0.41 53.00±0.00 49.27±0.09 50.27±0.09 51.20±0.33 52.80±0.28 53.80±0.16 0
BayesTune-LoRA 53.47±0.09 54.67±0.09 55.80±0.00 53.40±0.00 54.67±0.09 53.07±0.09 54.47±0.19 53.27±0.09 60.67±0.09 53.40±0.16 49.40±0.00 51.67±0.09 52.07±0.25 53.07±0.09 54.80±0.00 0
FedL2P 68.00±1.02 65.87±1.33 70.20±3.13 71.73±0.81 71.20±1.40 69.60±1.66 71.20±1.02 69.00±0.99 73.80±0.82 69.13±1.54 64.00±1.14 67.80±1.82 69.27±1.23 68.53±1.32 69.93±1.95 1
FedP2EFT 69.80±0.16 69.27±1.04 73.47±1.51 70.40±1.84 71.27±0.57 70.67±2.88 74.80±0.16 73.20±1.14 74.53±1.25 73.53±0.75 65.93±2.64 68.60±1.70 73.67±1.84 70.67±0.25 72.80±0.71 14
Table 5: Mean latency and memory costs across 100 runs of the first client in the seen pool using an NVIDIA A100 GPU (80GB).
Federated Training Inference
Dataset (Model) Approach Mean Latency (s) Peak Memory (GB) Mean Latency (s) Peak Memory (GB)
XNLI (mBERT) LoRA - - 3.16 2.69
AdaLoRA - - 3.41 2.77
BayesTune-LoRA - - 2.481 + 3.13 3.091 + 2.69
FedL2P 28.16 3.53 4.48 2.99
FedP2EFT 3.36 3.09 4.41 2.69
Fed-Aya (Llama-3.2-3B) FT - - 347.33 17.92
AdaLoRA - - 423.88 18.77
BayesTune-LoRA - - 72.291 + 357.56 25.181 + 18.33
FedL2P 226.49 32.19 379.98 19.78
FedP2EFT 80.28 25.19 400.64 15.67

### v2

@@ 4.3 Setup: Baselines
Given a base model, we compare FedP2EFT with existing fine-tuning and learning to personalize approaches. For each baseline, we either follow best practices recommended by the corresponding authors or employ a simple grid search and pick the best performing hyperparameters (Appendix A).
LoRA PEFT. We deploy LoRA (Hu et al. 2022) on all linear layers of the model with a fixed rank $r$.
Non-FL Rank Selection. We compare with AdaLoRA (Zhang et al. 2023) and our proposed LoRA-variant of BayesTune (Kim and Hospedales 2023), BT-LoRA (Section 3.2), which optimizes $\bm{\lambda}$ separately for each client.
FL to Personalize. We compare with FedL2P (Lee et al. 2023) which trains a MLP federatedly to output per-client learning rates for each LoRA module.
@@ 4.4 Results on Text Classification
We evaluate our approach in a typical FL setup, where the pretrained model is first trained using Standard FL with full fine-tuning and the resulting base model is then personalized to each client. Tables 1 & 2 show the mean and standard deviation (SD) of the accuracy for each language in our MasakhaNEWS setup for seen and unseen pool respectively (similarly for XNLI in Appendix Tables 13 & 14).
@@
The results in all four tables show that federated learning to personalize methods (FedL2P and FedP2EFT) outperform the other baselines in most cases. Non-FL rank selection approaches (AdaLoRA and BT-LoRA), on the other hand, tend to overfit and/or struggle to learn an optimal rank structure given the limited number of samples in each client. Comparing FedL2P and FedP2EFT, FedP2EFT largely surpass FedL2P with a few exceptions, indicating that learning to personalize LoRA rank structure is the better hyperparameter choice than personalizing learning rates; this finding is also aligned with recent LLM-based optimizer findings (Zhao et al. 2025), which shows that Adam’s performance is robust with respect to its learning rate.
FedP2EFT’s pFL Compatibility. Apart from Standard FL, we show that FedP2EFT can be plugged into existing pFL works that trains both a subset of the pretrained model and personalized layers for each client. Table 3 and Appendix Table 12 show that FedP2EFT outperforms baselines in almost all cases in our XNLI setup given a base model trained using FedDPA-T (Yang et al. 2024b) and DEPT(SPEC) (Iacob et al. 2025) respectively. In short, FedP2EFT can be integrated into a larger family of existing pFL approaches, listed in Section 2, to further improve personalization performance.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
bg 45.8±0.3 44.1±0.1 43.8±0.0 47.5±2.8 64.4±0.2
hi 42.8±0.2 41.2±0.0 40.6±0.0 44.5±3.2 57.8±1.3
es 48.7±0.2 47.5±0.1 47.2±0.0 50.3±2.9 58.5±1.3
el 50.9±0.1 50.0±0.0 50.0±0.0 51.5±1.4 59.7±2.6
vi 53.0±0.0 52.4±0.0 52.4±0.0 53.6±0.7 60.8±0.7
tr 48.0±0.3 46.5±0.1 46.2±0.0 50.1±2.8 58.9±2.9
de 49.9±0.1 48.0±0.0 47.4±0.2 50.9±2.6 55.0±0.2
ur 41.9±0.1 38.8±0.0 38.8±0.0 44.8±4.5 63.7±0.2
en 46.5±0.3 44.7±0.1 44.3±0.1 49.1±4.3 55.9±0.2
zh 44.4±0.2 42.2±0.0 41.6±0.0 46.5±4.0 56.3±0.3
th 42.5±0.2 40.7±0.1 40.5±0.1 44.4±3.7 58.3±0.2
sw 51.8±0.2 50.4±0.0 49.8±0.0 52.3±1.3 59.5±0.5
ar 46.9±0.2 45.0±0.0 44.7±0.1 48.4±3.0 55.2±0.2
fr 48.1±0.2 46.0±0.0 46.0±0.0 49.5±3.0 57.5±0.2
ru 50.5±0.2 48.8±0.0 48.1±0.2 51.6±2.3 55.2±0.4
Table 3: Mean±SD Accuracy of each language across 3 different seeds for clients in the seen pool of our XNLI setup ($r=2$, see Appendix Table 9 for other $r$ values). The pretrained model is trained using FedDPA-T and the resulting base model is personalized to each client given a baseline approach. See Appendix Table 12 for results with DEPT (SPEC).
4.5 Results on Instruction-Tuning Generation
We evaluate our approach on the more challenging real-world multilingual benchmark, Fed-Aya. Tables 4 and 5 show the average METEOR (Banerjee and Lavie 2005)/ROUGE-L (Lin 2004) [^5] Due to limited space, we include ROUGE-1 in the Appendix. of each language given the off-the-shelf instruction finetuned Llama-3.2-3B (Llama-3.2-3B-Instruct) for seen and unseen clients respectively. Similarly, in Appendix Tables 15 and 16, we show the same tables given a pretrained MobileLLaMA-1.4B model trained using Standard FL with LoRA following the training recipe from FedLLM-Bench (Ye et al. 2024). These two models represent scenarios where the base model may or may not be trained using FL.
In all four tables, FedP2EFT outperforms baselines in most scenarios. We also observe that FedL2P mostly underperforms standard baselines, a phenomenon also observed for our XNLI setup when the base model is trained with FedDPA-T (Appendix Table 9). We hypothesize that the inner-loop optimization in FedL2P fail to reach a stationary point [^6] FedL2P relies on the implicit function theorem for hypergradient computation. due to the inherent task difficulty (Fed-Aya) or a less-performant base model, resulting in a sub-optimal hypergradient and downstream performance.
FedP2EFT Limitations. In some cases, FedP2EFT falls short, especially in the recall performance (ROUGE), such as Russian (ru) and French (fr) for unseen clients for both base models in most scenarios. These cases highlight two limitations of our approach:
i) Challenge with zero-shot transfer. ru is not seen by PSG during federated training; there are no ru samples, nor other languages similar to ru, in the seen pool, resulting in worse ru performance. Hence, we do not expect a similar outcome in datasets with a more diverse pool of clients.
ii) Predominant language dependence. None of the clients in the unseen pool have fr as a predominant language (Fig. 3). Because the number of predominant language samples heavily influences $\bm{\lambda}$, the performance for fr is worse. Instance-level $\bm{\lambda}$ generation, though a solution, is too costly, requiring one PSG forward pass per sample. This calls for novel efficient solutions that can better handle each client’s minority languages and is left as future work.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.24/0.13 0.24/0.14 0.24/0.14 0.23/0.13 0.24/0.14
ar 0.34/0.07 0.34/0.07 0.32/0.06 0.33/0.07 0.37/0.08
es 0.39/0.39 0.39/0.40 0.38/0.38 0.38/0.38 0.39/0.39
en 0.33/0.31 0.35/0.33 0.29/0.23 0.33/0.30 0.37/0.36
fr 0.29/0.30 0.29/0.29 0.29/0.29 0.29/0.30 0.34/0.33
zh 0.11/0.12 0.11/0.12 0.09/0.11 0.09/0.12 0.11/0.12
pt 0.38/0.41 0.38/0.41 0.37/0.39 0.38/0.40 0.40/0.42
Table 4: Avg. METEOR/ROUGE-L for seen clients in our Fed-Aya setup. ($r=2$, see Appendix Table 10 for other $r$ values) Base model is off-the-shelf Llama-3.2-3B-Instruct.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
te 0.16/0.08 0.16/0.11 0.17/0.09 0.16/0.11 0.17/0.07
ar 0.24/0.04 0.23/0.07 0.22/0.05 0.24/0.05 0.26/0.05
es 0.43/0.44 0.43/0.44 0.38/0.41 0.41/0.43 0.44/0.48
en 0.32/0.25 0.32/0.25 0.32/0.24 0.31/0.24 0.31/0.27
fr 0.55/0.67 0.55/0.67 0.40/0.33 0.40/0.33 0.22/0.22
zh 0.25/0.00 0.25/0.00 0.23/0.00 0.23/0.00 0.28/0.01
pt 0.36/0.40 0.33/0.40 0.31/0.36 0.34/0.39 0.35/0.41
ru 0.22/0.17 0.23/0.17 0.25/0.18 0.25/0.20 0.34/0.27
Table 5: Avg. METEOR/ROUGE-L for unseen clients in our Fed-Aya setup ($r=2$, see Appendix Table 11 for other $r$ values). Base model is off-the-shelf Llama-3.2-3B-Instruct.
4.6 Cost of FedP2EFT
Table 6 shows the mean latency and the peak memory usage across 100 runs on the first client in the seen pool for $r=16$ using an Nvidia A100 GPU. Non-FL baselines do not incur a federated training cost while FL approaches requires training the PSG. Comparing FedL2P and FedP2EFT, FedP2EFT does not require second-order optimization, resulting in better efficiency. Note that FedL2P needs to be run for every $r$ while FedP2EFT runs once for all targeted ranks.
During inference, FL-based approaches incur an additional forward pass of base model and the PSG compared to non-FL approaches. Memory-wise, FedP2EFT results in the smallest memory footprint for autoregressive generation as the PSG learns not to attach LoRA modules $\lambda_{l}=0$ on some layers, skipping matmul operations entirely. More details discussions on cost can be found in Appendix B.
Federated Training
Dataset
Approach
Mean
Latency (s)
Peak
Memory (GB)
XNLI
FedL2P 28.2 3.5
FedP2EFT 3.4 3.1
Fed-Aya
(Llama-3.2-3B)
FedL2P 226.5 32.2
FedP2EFT 80.3 25.2
Inference
Dataset
Approach
Mean
Latency (s)
Peak
Memory (GB)
XNLI
LoRA 3.1 2.6
AdaLoRA 3.4 2.7
BT-LoRA 2.41 + 3.1 3.11 + 2.7
FedL2P 4.5 3.0
FedP2EFT 4.4 2.7
Fed-Aya
(Llama-3.2-3B)
LoRA 347.3 17.9
AdaLoRA 423.9 18.8
BT-LoRA 72.31 + 357.6 25.21 + 18.3
FedL2P 380.0 19.8
FedP2EFT 400.6 15.7